differential equations

Use the Laplace transform to solve the given initial-value problem. y + 3y = et, y(0) = 2 y(t) =
Answer #1

Given eq? is y +34= est 4 (0)=2 taking laplace transform of each term 2 (4) +32(Y)=2(est) 1 2 face) s 9 SY(S)-4(0) + 3 Ycs)